A,B and C try to hit a target.The respective chances of them hitting the target are :for A,1/2 ,for B,2/3 ,for C 1/4.What is the probability that A or B hit the target?Please explain clearly.Is the answer 3/8?

Grade:12

1 Answers

vikas askiitian expert
13 years ago

p(A) =1/2 , P(A)1=1/2

p(B) =2/3 , p(B)1=1/3

p(C)=1/4 , p(C)1=3/4

 

possible cases => {(Ahits,Bnot,Cnot) ,(Ahits,Bnot,Chits) (Bhits,Anot,Cnot),(Bhits,Anot,Chits)}

adding all these =>{1/2*1/3*3/4 + 1/2*1/3*1/4 + 2/3*1/2*3/4 + 2/3*1/2*1/4}

                     =1/2